OpenStudy (anonymous):

f(x) = 4throot(x); find the slope and the y-intercept of the tangent line T at x = 1.

OpenStudy (anonymous):

f'(x) = 1/(4x^[3/4]), f'(1) = 1/4 = m, f(1) = 1, T - 1 = (1/4)(x-1), T = (1/4)(x) + 3/4

OpenStudy (anonymous):

right?